A 0.700 mm -diameter ball bearing has 2.10×109 excess electrons. What is the ball bearing's potential?
Answer #1

e = magnitude of charge on electron = 1.6 x 10-19 C

n = number of excess electrons = 2.10 x 109 electrons

Q = magnitude of charge

Magnitude of charge is given as

Q = ne

Q = (2.10 x 109) (1.6 x 10-19)

Q = 3.36 x 10-10 C

d = diameter of the ball = 0.700 mm = 0.7 x 10-3 m

r = radius of the ball = d/2 = (0.7 x 10-3 )/2 = 3.5 x 10-4 m

Ball bearing's potential is given as

V = kQ/r

V = (9 x 109) (3.36 x 10-10)/(3.5 x 10-4 )

V = 8640 Volts